OpenStudy (anonymous):

Determine whether the point at (-2, -16) is the location of a minimum or maximum, or a point of inflection for the function x^2 +4x - 12

OpenStudy (anonymous):

find the derivative of the function , the derivative =0 then find x

OpenStudy (anonymous):

whats the derivative of the function?

OpenStudy (anonymous):

\[f \prime(x) =2x+4\]\[f \prime(x)=0\] then 2x+4=0 divide by 2 x+2=0 x=-2

OpenStudy (anonymous):

u must know the derivative rules to find the derivative

OpenStudy (anonymous):

ohhhh thank you!

OpenStudy (anonymous):

the solution is not over yet

OpenStudy (anonymous):

u must find the the second derivative \[f \prime \prime (x)=2 \] \[f \prime \prime (-2)=2 \] 2>0 then the point (-2,16) is minimum

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

and u r welcome :D

OpenStudy (anonymous):

ohh okay yes i get it, thank you again.

OpenStudy (anonymous):

u r welcome again :D , anytime